Given that the complex number z and its conjugate z* satisfy the equation zz* + 2iz = 12 +6i. Find the possible value of z.



1
Expert's answer
2021-09-06T14:27:49-0400

"z=x+iy"

"z^*=x-iy"

"zz^*=x^2+y^2"


"x^2+y^2+2xi-2y=12+6i"

"x^2+y^2-2y=12"

"2xi=6i"


"x=3"

"9+y^2-2y=12"

"y^2-2y-3=0"


"y_1=\frac{2-\sqrt{4+12}}{2}=-1"


"y_2=3"


Answer:


"z_1=3-i"

"z_2=3+3i"